State Farm Insurance studies show that in Colorado, 45% of the auto insurance claims submitted for property damage were submitted by males under 25 years of age. Suppose 12 property damage claims involving automobiles are selected at random.

(a) Let r be the number of claims made by males under age 25. Make a histogram for the r-distribution probabilities.

Maple Generated Plot Maple Generated Plot
Maple Generated Plot Maple Generated Plot



(b) What is the probability that three or more claims are made by males under age 25? (Use 3 decimal places.)


(c) What is the expected number of claims made by males under age 25? What is the standard deviation of the r-probability distribution? (Use 2 decimal places.)

μ
σ
Answer #1

Part a)

B :- Binomial Distribution

X ~ B ( n = 12 , P = 0.45 )
P(X = 2) = CHPQin-r)
C = n! (n - 1)!?!

X P(X)
0 0.0008
1 0.0075
2 0.0339
3 0.0923
4 0.17
5 0.2225
6 0.2124
7 0.1489
8 0.0762
9 0.0277
10 0.0068
11 0.001
12 0.0001

Histogram 1 2 3 4 5 6 7 8 9 10 11 12 13

0.15 plo 0 0.05 0 1 2 3 4 5 6 7 8 9 10 11 12

Part b)

P ( X >= 3 ) = 1 - P ( X <= 2 )

P(X = 0) = C12(0.45)º(0.55) (12-0) = 0.0008
P(X = 1) = C120.45)*(0.55) (12-1) = 0.0075
P(X = 2) = 0.120.45)2(0.55) (12-2) = 0.0339

P ( X >= 3 ) = 1 - 0.0422 = 0.958

Part c)

Mean \mu = n * P = ( 12 * 0.45 ) = 5.4
Variance \sigma^2 = n * P * Q = ( 12 * 0.45 * 0.55 ) = 2.97
Standard deviation \sigma    = (variance) = (2.97) = 1.72